Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name jmammei
Log entry time 16:00:13 on October 10, 2010
Entry number 334238
keyword=Day shift summary
Summary for October 10, Day Shift
People in counting house:
- SL: Juliette Mammei
- TO: Vincenzo Bellini
- 3rd: Shloka Chandavar
- Others: Sasha, Carlos, Alex, Jixie, Roman
Main Events/Issues:
- 08:00 Beam down due to ARC dipole trip
- 08:10 L_MAG alarm - clears itself before I can see what it is
- 08:30 MCC estimates ~15 minutes until sending tune beam
- 08:45 L_MAG alarms 2 times in quick succession - second time,
lock lost but "fixed itself" after about 30 seconds
- 10:00 MCC starting to try to send tune beam in the injector
- 10:26 moved the target out for Moller measurements
- 11:14 turned off spectrometer HV for cosmic testing with the
calorimeter
- 11:30 cryo temp dropped b/c of issue with Hall C target
- 12:00 Bit of a bottleneck with ARM/operators - Hall C is
using the ARM.
- 12:10 Hall C finished with the survey - sent the ARM to Hall
A - Jixie and Carlos go in.
- 12:46 Jixie and Carlos out of hall
- 14:02 Called for a controlled access for Moller: Sasha in and
out.
- 14:30 Turned on spectrometer HV in preparation for resuming
the run plan.
- 15:00 Get beam back; perform beam recovery procedure - raster
increased to 4x4mm^2
- 15:25 Move to 15 cm LH2 target, do spot++
- 15:45 Begin the first elastic calibration run of this shift
- 16:00 End shift with elastic calibration in progress
Summary:
- There were multiple alarms for the NMR and it lost the lock,
though it seems to recover by itself. There was no beam all morning.
Sasha and crew came in and there was still no beam, but MCC was sending
tune beam to the injector. Miscommunication caused us to think there was
a bottleneck with the ARM - MCC didn't mention that Hall C wanted a
survey (and not a long escorted access) which caused some consternation
on the part of the Moller folks who needed to fill the Moller magnet.
- There was some difficulty in turning on the solenoid, so the
Moller running was postponed and we prepared for resuming the run plan.
- Resumed the run plan at the lowest momentum setting (-4%) for
the seive-out elastic calibration. Need 3 additional 1-hour long runs
for this setting.
Run Summary:
- 7111-7120 - Tests; Alex did some cosmics, DAQ tests and
- 7121 - Empty target, 5uA, spot++ run (rates too low)
- 7122 - Empty target, 5uA, spot++ run - looks good
- 7123 - 15cm LH2 target, 3uA, spot++ run
- 7124 - 15cm LH2 target, 10uA, 3.194 GeV/c (p0=-4%), sieve out
- first calibration run this shift
A copy of this log entry has been emailed to: riordan@jlab.org, munoz@jlab.org, jroche@jlab.org, camsonne@jlab.org, sabatie@jlab.org